A dog toy is thrown from 5 inches above the ground. After 2 seconds, the toy reaches a maximum height of 9 inches, and then lands back on the ground 5 seconds after it was thrown. Find the "a" (coefficient of the x^2) and write this in quadratic equation in vertex form, standard form, and factored form.
Answer by stanbon(75887)

A dog toy is thrown from 5 inches above the ground. After 2 seconds, the toy reaches a maximum height of 9 inches, and then lands back on the ground 5 seconds after it was thrown. Find the "a" (coefficient of the x^2) and write this in quadratic equation in vertex form, standard form, and factored form.
========
Form:: h(t) = ax^2 + bx + c
-----
You have points (0,5),(2,9),(5,0)
-----
Using (0,5) you get: 0 + 0 + c = 5
Using (2,9) you get: 4a+2b + c = 9
Using (5,0) you get: 25a+5b+ c = 0
------------------------------------------
Since c = 5 you get:
4a + 2b = 4
25a +5b = -5
---------
Modify for elimination:
2a + b = 2
5a + b = -1
------
3a = -3
a = -1
----
b = 2-2a = 2+2 = 4
c = 5
---
Ans: h(x) = -x^2 + 4x + 5
======
Cheers,
Stan H.
-----------